80-310/610 Logic and Computation Fall 2001

Homework #1

Due Wednesday, August 29

1. Take the self-test but do not turn it in. Talk to me if you are not sure whether or not this course is right for you.

2. Read the introduction to the van Dalen text.

3. Read section 1.2 from Enderton, A Mathematical Introduction to Logic, at the end of the course notes.